OpenStudy (keeli_grace):

For the function f(x) = 2 – 3x, find f(–4). A. –5 B. –10 C. 14 D. 12

OpenStudy (anonymous):

just sub in -4 for every x f(x) = 2 - 3x f(-4) = 2 - 3(-4) f(-4) = 2 + 12 f(-4) = 14 understand ?

OpenStudy (anonymous):

To find f(-4) just input 'x' for -4. f(x) = 2 - 3x f(x) = 2 - 3(-4) f(x) = 2 + 12 f(x) = 14

OpenStudy (anonymous):

Lol beat me to it @kelliegirl33

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

Oops. To find f(-4) just input '-4' for 'x'.*

OpenStudy (anonymous):

:p

OpenStudy (keeli_grace):

Thank you @kelliegirl33 & @greenlegodude57 :)